### v3

# End-to-End Energy Saving in Cell-Free Massive MIMO ISAC for Ultra-Reliable Target-Aware Actuation

Abstract
[…]
Index Terms:
Integrated sensing and communication, cell-free massive MIMO, URLLC, power allocation, blocklength
I Introduction
6G mobile networks are expected to offer various sensing-based applications such as autonomous vehicles, smart homes/cities/factories, remote healthcare, industrial Internet-of-things (IIoT), and robot control in target/environment-aware scenarios [1, 2]. These applications introduce ultra-reliable target-aware actuation use cases, which involve scenarios where a communication and sensing system must detect the presence, position, and motion of a target (e.g., a human or object) with high accuracy and minimal delay, and then trigger an actuation response (e.g., controlling a robot, vehicle, or machine) with ultra-reliability and low latency.
To support such use cases, the system must deliver sensing information to user equipment (UE) with a minimum reliability of 99.999% and an end-to-end (E2E) latency of less than 10–150 ms [3, 4, 5]. This demand necessitates two key aspects defined in 6G networks: integrated sensing and communication (ISAC) together with ultra-reliable low-latency communication (URLLC).
In ultra-reliable target-aware actuation use cases such as traffic control and autonomous vehicles, real-time transmission of sensing information to URLLC UEs is essential, as the timeliness of updates directly affects the performance and safety. To capture the frequency at which sensing information must be updated, this paper introduces the concept of the “refreshing rate”, defined by 3GPP [6] as the number of sensing measurements and updates delivered per unit time [7]. The refreshing rate is closely tied to the communication blocklength: longer blocklengths lead to lower refreshing rates due to the increased time required for each sensing operation. This trade-off highlights the potential for optimizing blocklength to achieve a balance between reliable communication and a high refreshing rate. Although URLLC and ISAC have typically been studied as separate domains, a joint design and optimization approach is required to meet the distinct requirements of both and maximize integration gains.
Cell-free massive multiple-input multiple-output (CF-mMIMO) has emerged as a strong candidate for meeting the stringent requirements of URLLC applications [8, 9, 10]. It offers high reliability by compensating for large path-loss variations and significantly improves the performance of cell-edge UEs. Moreover, CF-mMIMO is well-suited for implementing ISAC, as its distributed access points (APs) can support bi-static and multi-static sensing configurations [11, 12, 13, 14]. These configurations eliminate the need for full-duplex capabilities at individual APs but require effective coordination among them. Centralized radio access network (C-RAN) architectures, which support centralized processing of network functions, enables joint transmission/reception and processing [15]. This centralized approach is also essential for facilitating bi-static and multi-static sensing.
While CF-mMIMO networks offer significant performance advantages, the integration of sensing functionalities is anticipated to substantially increase transmission energy consumption, resulting in up to a tenfold reduction in energy efficiency [16]. This challenge is further compounded by the fact that the combined demands of baseband processing and sensing operations may result in higher processing energy consumption compared to conventional communication-only systems. Most existing research on green CF-mMIMO without sensing integration has primarily focused on optimizing transmit power, with relatively limited attention given to processing energy consumption, aside from a few notable exceptions [17, 18]. Moreover, the impact of sensing functionalities on processing energy consumption in such networks remains insufficiently explored.
It is important to mention that communication system efficiency is typically measured in terms of data rate while sensing tasks are evaluated by task-specific metrics such as target detection success rates. The execution of a sensing task involves not only signal transmission and reception of reflections from the target but also extensive processing to extract the desired information. This post-reception processing is a fundamental component of completing a sensing task, and thus must be explicitly accounted for in system design and analysis. This distinction highlights the need for a comprehensive E2E energy consumption analysis that accounts for both transmission and computation costs.
The virtualization capabilities offered by C-RAN architectures facilitate centralized resource management and flexible allocation of processing resources across APs and tasks. This architectural flexibility presents a promising opportunity for E2E optimization of both processing and radio resources in ISAC networks.
Motivated by these considerations, this paper addresses the following key research question: How can sensing be effectively integrated into CF-mMIMO systems with CRAN architecture and supporting URLLC, while minimizing E2E energy consumption?
I-A Literature Review
The finite blocklength regime and URLLC have primarily been studied in the context of cellular networks (see [19, 20, 21] and references therein), with limited research addressing URLLC in CF-mMIMO networks. The potentials of massive MIMO in meeting URLLC requirements for industrial automation are investigated in [22, 23, 24]. Network availability is investigated in [22] and [23], and [24] proposes a power allocation algorithm to maximize the minimum signal-to-interference-plus-noise ratio (SINR). The authors in [25] propose two power optimization algorithms to provide URLLC for traditional ground UEs and unmanned aerial vehicles (UAVs) in a CF-mMIMO system. The objectives in [25] are maximizing either the sum or the minimum URLLC rate, and they study neither energy minimization nor blocklength optimization. Max-min rate optimization is also studied in [10], where the authors adapt a special class of conjugate beamforming for a CF-mMIMO with single-antenna APs. Moreover, the authors aim to address an energy efficiency optimization problem.
Joint power and blocklength optimization is studied in [26] and [27] to optimize the reliability in a factory automation scenario and heterogeneous cellular networks, respectively. Cell-free networks as well as energy minimization are not studied in these works. Moreover, sensing requirements are not considered in [24, 22, 25, 10, 26, 27, 23].
There are few works that jointly consider URLLC and ISAC. In [5], a joint precoding scheme is proposed to minimize transmit power, satisfying sensing and delay requirements. Moreover, joint ISAC beamforming and scheduling design is addressed in [28] and [29] with a focus on the coexistence of periodic and aperiodic traffic to balance the tradeoff between their corresponding performance. The aperiodic traffic is triggered by sensing information about the stochastic environment.
The consideration of E2E energy-awareness has been explored in various contexts, as reflected in prior works such as [30, 31, 17, 18, 32]. In particular, [18] studied fully virtualized E2E power minimization problem for CF-mMIMO on O-RAN architecture by taking the radio, fronthaul, and processing resources into account. Joint UE scheduling and power allocation scheme for cell-free ISAC systems is studied in [33], which aims to maximize the sum rate of the communication UEs and the sensing targets. However, the authors do not consider either URLLC requirements or energy minimization.
TABLE I: NOVELTY COMPARISON OF OUR PAPER TO THE EXISTING WORKS IN LITERATURE
[19, 20, 21] [22, 23, 24] [25] [10] [26, 27] [5] [28, 29] [30, 31, 17, 18, 32] [16] This paper
ISAC
URLLC
MIMO
Cell-free networks
Multi-static sensing
Target detection
E2E energy consumption
Blocklength optimization
Table I presents a comparison highlighting the novelty of this work relative to existing studies. To the best of our knowledge, our previous work [16] was the first to investigate ISAC in CF-mMIMO systems with URLLC UEs, with a specific focus on energy efficiency. However, the impact of blocklength optimization was not addressed in that study. In this paper, we formulate and solve a joint blocklength and power optimization problem from an E2E energy minimization perspective. E2E energy consumption encompasses energy used both at the distributed radio APs and at the centralized cloud infrastructure.
Figure 1: ISAC network architecture in CF-mMIMO with URLLC UEs and a joint resource allocation mechanism.
I-B Contributions
Integration of communication and sensing functionalities in a CF-mMIMO network is depicted in Fig. 1. The objective is to jointly optimize blocklength and power allocation to minimize E2E energy consumption across both the radio access and edge cloud domains. URLLC requirements—such as reliability and latency—and sensing requirements—such as refreshing rate and accuracy—are jointly addressed through coordinated resource allocation. The edge cloud manages centralized processing for both sensing and communication, while distributed APs enable multi-static sensing alongside communication services. This architecture leverages shared radio and computation resources to facilitate joint performance optimization across sensing and communication domains.
In this paper, we study E2E energy consumption in a CF-mMIMO system with URLLC UEs and multi-static sensing in a cluttered environment. We sense a candidate location of the potential target utilizing communication signals and an additional sensing signal. Since the sensing signals can cause interference for the UEs, we employ null-space precoding vector for the sensing signal to null the interference to the UEs. The main contributions of this paper are outlined as follows:
  • We conduct a comprehensive analysis of the processing requirements, specifically in terms of giga-operations per second (GOPS) for communication and sensing tasks at the cloud, and provide analytical expression of the cloud energy consumption.
  • We present an E2E energy minimization algorithm by jointly optimizing the blocklength and power and compare it with two baseline algorithms. The first baseline algorithm aims to minimize the E2E energy consumption for only CF-mMIMO systems with URLLC without sensing capability. The second baseline considers sensing in the system while it aims to minimize only transmission energy consumption at the radio sites. To address the non-convex optimization challenges, we employ feasible point pursuit - successive convex approximation (FPP-SCA), concave-convex programming (CCP), and fractional programming.
  • We evaluate two distinct target detection schemes based on the maximum a posteriori ratio test (MAPRT), each with different processing requirements—namely, the clutter-unaware detector and the clutter-aware detector. Their performance is compared in terms of E2E energy consumption and detection probability.
  • We conduct sensitivity analyses to evaluate the effects of communication and sensing requirements, number of antennas per AP, and sensing receive APs on E2E energy consumption and sensing performance.
The rest of the paper is organized as follows: Section II introduces the system model. Section III provides URLLC analysis, considering decoding error probability (DEP) and delay in the finite blocklength regime. Section IV covers the sensing analysis. Section V describes the E2E power model and derives GOPS analysis for both communication and sensing. Optimization problems are presented in Section VI, followed by numerical results and conclusions in Sections VII and VIII, respectively.
[…]
II System Model
We study ISAC in a CF-mMIMO system in URLLC scenarios. The system adopts a virtualized C-RAN architecture [34] for uplink channel estimation and downlink communication, as well as multi-static sensing as shown in Fig. 2. All the APs are interconnected via fronthaul links to the edge cloud and operate in full synchronization. Notice that synchronization presents a major challenge in multi-static sensing, as asynchronous transmitters and receivers can lead to ambiguity and performance degradation. While this issue has been studied in the context of communication in cell-free networks [35], the synchronization mechanisms already developed for CF-mMIMO communication make it a promising infrastructure for implementing advanced sensing techniques. This motivates our focus on exploring other critical aspects of CF-mMIMO-based ISAC systems. Another important consideration is the significant fronthaul capacity requirement imposed by centralized signal processing for both communication and sensing functions [36]. While this paper adopts a centralized processing framework to evaluate the full potential of CF-mMIMO ISAC systems, addressing fronthaul limitations remains an open research direction.
We consider the original form of CF-mMIMO [8], wherein all the $N_{\rm tx}$ ISAC APs jointly serve the $N_{\rm ue}$ URLLC UEs by transmitting precoded signals containing both communication and sensing symbols.
Figure 2: ISAC system model in CF-mMIMO with URLLC UEs.
Concurrently, the $N_{\rm rx}$ sensing receive APs engage in simultaneous sensing of the candidate location to detect the target. Each AP is equipped with an array of $M$ antennas configured in a horizontal uniform linear array (ULA) with half-wavelength spacing. The respective array response vector is $\textbf{a}(\varphi,\vartheta)=\begin{bmatrix}1\!&\!e^{j\pi\sin(\varphi)\cos(\vartheta)}&\ldots&e^{j(M-1)\pi\sin(\varphi)\cos(\vartheta)}\end{bmatrix}^{T},$ where $\varphi$ and $\vartheta$ are the azimuth and elevation angles from the AP to the target location, respectively [37].
We consider the finite blocklength regime for URLLC UEs, where a packet of $b_{i}$ bits is sent to UE $i$ within a transmission block with blocklength $L=L_{p}+L_{d}$ symbols using the coherence bandwidth $B$. $L_{p}$ and $L_{d}$ are the number of symbols for pilot and data, respectively. It is expected that duration of each URLLC transmission, denoted by $T$, is shorter than one coherence time $T_{c}$, i.e., $T<T_{c}$ [10]. Without loss of generality, we assume that we have a channel estimation phase in each transmission.
II-A Downlink ISAC Transmission
Let $s_{i}[m]$ and $s_{0}[m]$ represent the downlink communication symbol for UE $i$ and sensing symbol, respectively at time instance $m$. The symbols are independent and have zero mean and unit power. Moreover, let $\rho_{i}\geq 0$ and $\rho_{0}\geq 0$ be, respectively, the power control coefficients for UE $i$ and the target, which are fixed throughout the transmission. Then, the transmitted signal from transmit AP $k\in\{1,\ldots,N_{\rm tx}\}$ at time instance $m$ is
$\vskip-2.84526pt\textbf{x}_{k}[m]=\sum_{i=0}^{N_{\rm ue}}\sqrt{\rho_{i}}\textbf{w}_{i,k}s_{i}[m]=\textbf{W}_{k}\textbf{D}_{\rm s}[m]\bm{\rho},\vskip-2.84526pt$ (1)
where the vectors $\textbf{w}_{i,k}\in\mathbb{C}^{M}$ and $\textbf{w}_{0,k}\in\mathbb{C}^{M}$ are the transmit precoding vectors for transmit AP $k$ corresponding to UE $i$ and the sensing signal, respectively. In (1), $\textbf{W}_{k}=\begin{bmatrix}\textbf{w}_{0,k}&\textbf{w}_{1,k}&\cdots&\textbf{w}_{N_{\rm ue},k}\end{bmatrix}$, ${\textbf{D}}_{\rm s}[m]=\textrm{diag}\left(s_{0}[m],s_{1}[m],\ldots,s_{N_{\rm ue}}[m]\right)$ is the diagonal matrix containing the sensing and communication symbols, and $\bm{\rho}=[\sqrt{\rho_{0}}\ \ldots\ \sqrt{\rho_{N_{\rm ue}}}]^{T}$.
The communication channels are modeled as spatially correlated Rician fading, which are assumed to remain constant during each coherence block, and the channel realizations are independent of each other. Let $\textbf{h}_{i,k}\in\mathbb{C}^{M}$ denote the channel between ISAC AP $k$ and UE $i$, modeled as
$\textbf{h}_{i,k}=e^{j\varphi_{i,k}}\bar{\textbf{h}}_{i,k}\,+\,\tilde{\textbf{h}}_{i,k},$ (2)
which consists of a semi-deterministic line-of-sight (LOS) path, represented by $e^{j\varphi_{i,k}}\bar{\textbf{h}}_{i,k}$ with unknown phase-shift $\varphi_{i,k}\sim\mathcal{U}[0,2\pi)$, i.e., uniformly distributed on $[0,2\pi)$, and a stochastic non-LOS (NLOS) component $\tilde{\textbf{h}}_{i,k}\sim\mathcal{CN}(\textbf{0},\textbf{R}_{i,k})$ with the spatial correlation matrix $\textbf{R}_{i,k}\in\mathbb{C}^{M\times M}$. Both $\bar{\textbf{h}}_{i,k}$ and $\textbf{R}_{i,k}$ include the combined effect of geometric path loss and shadowing. We concatenate the channel vectors $\textbf{h}_{i,k}$ in the collective channel vector
$\textbf{h}_{i}=\begin{bmatrix}\textbf{h}_{i,1}^{T}&\ldots&\textbf{h}_{i,N_{\rm tx}}^{T}\end{bmatrix}^{T}\in\mathbb{C}^{N_{\rm tx}M},$ (3)
for UE $i$. The received signal at UE $i$ is given as
$y_{i}[m]=$ $\underbrace{\sqrt{\rho_{i}}\textbf{h}_{i}^{H}\textbf{w}_{i}s_{i}[m]}_{\textrm{Desired signal}}+\underbrace{\sum_{j=1,j\neq i}^{N_{\rm ue}}\sqrt{\rho_{j}}\textbf{h}_{i}^{H}\textbf{w}_{j}s_{j}[m]}_{\textrm{Interference signal due to the other UEs}}$
$+\underbrace{\sqrt{\rho_{0}}\textbf{h}_{i}^{H}\textbf{w}_{0}s_{0}[m]}_{\textrm{Interference signal due to the sensing}}+\underbrace{n_{i}[m]}_{\textrm{Noise}},$ (4)
where $n_{i}[m]\!\sim\!\mathcal{CN}(0,\sigma_{n}^{2})$ is the independent receiver noise at UE $i$ at time instance $m$ and the collective precoding vectors $\textbf{w}_{i}=\left[\textbf{w}_{i,1}^{T}\ \textbf{w}_{i,2}^{T}\ \ldots\ \textbf{w}_{i,N_{\rm tx}}^{T}\right]^{T}\in\mathbb{C}^{N_{\rm tx}M},$ for $i=1,\ldots,N_{\rm ue}$, and $\textbf{w}_{0}=\left[\textbf{w}_{0,1}^{T}\ \textbf{w}_{0,2}^{T}\ \ldots\ \textbf{w}_{0,N_{\rm tx}}^{T}\right]^{T}\in\mathbb{C}^{N_{\rm tx}M}$ are the centralized precoding vectors.
The communication and sensing transmit precoding vectors are obtained based on regularized zero forcing (RZF) and zero forcing (ZF) approaches, respectively. The unit-norm RZF precoding vector for UE $i$ is given as $\textbf{w}_{i}=\frac{\bar{\textbf{w}}_{i}}{\left\|\bar{\textbf{w}}_{i}\right\|}$, with
$\bar{\textbf{w}}_{i}\!=\!\left(\sum\limits_{j=1}^{N_{\rm ue}}\hat{\textbf{h}}_{j}\hat{\textbf{h}}_{j}^{H}+\delta\textbf{I}_{N_{\rm tx}M}\right)^{\!\!-1}\!\!\hat{\textbf{h}}_{i},\quad i=1,\ldots,N_{\rm ue},$ (5)
where $\delta$ is the regularization parameter, and $\hat{\textbf{h}}_{j}=\begin{bmatrix}\hat{\textbf{h}}_{j,1}^{T}&\ldots&\hat{\textbf{h}}_{j,N_{\rm tx}}^{T}\end{bmatrix}^{T}\in\mathbb{C}^{N_{\rm tx}M}$ is the linear minimum mean-squared error (LMMSE) channel estimate of the communication channel $\textbf{h}_{j}$, obtained as in [38]111 We omit the explanations from this paper due to the space limitation.. If the number of UEs is larger than the number of mutually orthogonal pilot sequences, then each pilot sequence may be assigned to multiple UEs using the pilot assignment algorithm in [34, Algorithm 4.1].
We aim to null the destructive interference from the sensing signal to the UEs by using the unit-norm ZF sensing precoding vector $\textbf{w}_{0}=\frac{\bar{\textbf{w}}_{0}}{\left\|\bar{\textbf{w}}_{0}\right\|}$, where
$\bar{\textbf{w}}_{0}=\left(\textbf{I}_{N_{\rm tx}M}-\textbf{U}\textbf{U}^{H}\right)\textbf{h}_{0},$ (6)
and U is the unitary matrix with the orthogonal columns that span the column space of the matrix $\begin{bmatrix}\hat{\textbf{h}}_{1}&\ldots&\hat{\textbf{h}}_{N_{\rm ue}}\end{bmatrix}$. $\textbf{h}_{0}=\begin{bmatrix}\sqrt{\beta_{1}}\textbf{a}^{T}(\varphi_{1},\vartheta_{1})&\ldots&\sqrt{\beta_{N_{\rm tx}}}\textbf{a}^{T}(\varphi_{N_{\rm tx}},\vartheta_{N_{\rm tx}})\end{bmatrix}^{T}\in\mathbb{C}^{N_{\rm tx}M}$ is the concatenated sensing channel between all the ISAC APs and the target, including the corresponding channel gains $\beta_{k}$ and the array response vectors $\textbf{a}(\varphi_{k},\vartheta_{k})$ for $k=1,\ldots,N_{\rm tx}$.
II-B Multi-Static Sensing
We employ multi-static sensing, utilizing multiple transmit and receive APs within the network. Target sensing is performed during the downlink phase. A LOS connection is assumed to exist between the target location and each participating transmit and receive AP. In the presence of the target, each receive AP captures both the reflected signals from the target and undesired signals, referred to as clutter. The clutter, being independent of the target’s presence, is treated as interference for sensing purposes. Without loss of generality, we assume that the LOS paths between transmit and receive APs are known and can be effectively canceled out. Consequently, the interference signals correspond to the reflected paths through obstacles and are henceforth denoted as target-free channels.
Let $\textbf{H}_{r,k}\in\mathbb{C}^{M\times M}$ denote the target-free channel matrix between transmit AP $k$ and receive AP $r$, which cause interference for sensing. We use the correlated Rayleigh fading model for the NLOS channels $\textbf{H}_{r,k}$, which is written as $\textbf{H}_{r,k}=\textbf{R}^{\frac{1}{2}}_{{\rm rx},(r,k)}\textbf{W}_{{\rm ch},{(r,k)}}\left(\textbf{R}^{\frac{1}{2}}_{{\rm tx},(r,k)}\right)^{T}$ using the Kronecker model [39]. The matrix $\textbf{W}_{{\rm ch},(r,k)}\in\mathbb{C}^{M\times M}$ is a random matrix with independent and identically distributed (i.i.d.) entries with $\mathcal{CN}(0,1)$ distribution. The matrix $\textbf{R}_{{\rm rx},(r,k)}\in\mathbb{C}^{M\times M}$ represents the spatial correlation matrix corresponding to receive AP $r$ and with respect to the direction of transmit AP $k$. Similarly, $\textbf{R}_{{\rm tx},(r,k)}\in\mathbb{C}^{M\times M}$ is the spatial correlation matrix corresponding to transmit AP $k$ and with respect to the direction of receive AP $r$. The channel gain is determined by the geometric path loss and shadowing, and is included in the spatial correlation matrices. The received signal at AP $r$ in the presence of the target and for $m\in\{1,\ldots,L_{d}\}$, can be formulated as
$\textbf{y}_{r}[m]=\sum_{k=1}^{N_{\rm tx}}\underbrace{\alpha_{r,k}\sqrt{\beta_{r,k}}\textbf{a}(\phi_{r},\theta_{r})\textbf{a}^{T}(\varphi_{k},\vartheta_{k})\textbf{x}_{k}[m]}_{\textrm{desired reflections from the target}}$
$+\sum_{k=1}^{N_{\rm tx}}\underbrace{\textbf{H}_{r,k}\textbf{x}_{k}[m]}_{\textrm{clutter}}+\textbf{n}_{r}[m],$ (7)
where $\textbf{n}_{r}[m]\sim\mathcal{CN}(\textbf{0},\sigma_{n}^{2}\textbf{I}_{M})$ is the receiver noise at the $M$ antennas of receive AP $r$. The second term in (II-B) refers to clutter which acts as interference for the target detection. Here, $\beta_{r,k}$ is the channel gain, including the path loss from transmit AP $k$ to receive AP $r$ through the target and the variance of bi-static radar cross-section (RCS) of the target denoted by $\sigma_{\rm rcs}$. $\beta_{r,k}$ is computed according to the radar range equation for bi-static sensing in [7, Chap. 2], given by
$\beta_{r,k}=\frac{\lambda_{c}^{2}\,\sigma_{\rm rcs}}{(4\pi)^{3}d_{{\rm tx},k}^{2}d_{{\rm rx},r}^{2}}.$ (8)
[…]
Following the same notation as in [40], in (II-B), the known part of each reflected path is denoted by ${\textbf{g}_{r,k}[m]\in\mathbb{C}^{M}}$, defined as
$\textbf{g}_{r,k}[m]\triangleq\sqrt{\beta_{r,k}}\textbf{a}(\phi_{r},\theta_{r})\textbf{a}^{T}(\varphi_{k},\vartheta_{k})\textbf{x}_{k}[m],$ (9)
where the matrix $\alpha_{r,k}\sqrt{\beta_{r,k}}\textbf{a}(\phi_{r},\theta_{r})\textbf{a}^{T}(\varphi_{k},\vartheta_{k})$ represents the reflected path through the target. Here, $\phi_{r}$ and $\theta_{r}$ denote the azimuth and elevation angles from the target location to receiver AP $r$, respectively. Similarly, $\varphi_{k}$ and $\vartheta_{k}$ represent the azimuth and elevation angles from transmit AP $k$ to the target location.
Each receive AP sends their respective signals $\textbf{y}_{r}[m]$, for $r=1,\ldots,N_{\rm rx}$, to the edge cloud to form the collective received signal $\textbf{y}[m]=\begin{bmatrix}\textbf{y}_{1}^{T}[m]&\ldots&\textbf{y}_{N_{\rm rx}}^{T}[m]\end{bmatrix}^{T}$, which can be expressed as
$\textbf{y}[m]$ $=\underbrace{\mathrm{blkdiag}\left(\textbf{G}_{1}[m],\ldots,\textbf{G}_{N_{\rm rx}}[m]\right)}_{\triangleq\textbf{G}[m]}\bm{\alpha}$
$+\underbrace{\left(\textbf{I}_{N_{\rm rx}}\otimes\left(\textbf{x}^{T}[m]\otimes\textbf{I}_{M}\right)\right)}_{\triangleq\textbf{X}[m]}\bm{\mathfrak{h}}+\textbf{n}[m]$
$=\textbf{G}[m]\bm{\alpha}+\textbf{X}[m]\bm{\mathfrak{h}}+\textbf{n}[m],$ (10)
where $\textbf{G}_{r}[m]=\begin{bmatrix}\textbf{g}_{r,1}[m]&\ldots&\textbf{g}_{r,N_{\rm tx}}[m]\end{bmatrix}\in\mathbb{C}^{M\times N_{\rm tx}}$, $\textbf{x}[m]=\begin{bmatrix}\textbf{x}^{T}_{1}[m]&\ldots&\textbf{x}^{T}_{N_{\rm tx}}[m]\end{bmatrix}^{T}\in\mathbb{C}^{N_{\rm tx}M}$, and $\bm{\mathfrak{h}}\sim\mathcal{CN}\left(\textbf{0},\textbf{R}\right)$ is the vectorized target-free channel[40, Sec. V]. In addition, we have defined $\textbf{G}[m]=\mathrm{blkdiag}\left(\textbf{G}_{1}[m],\ldots,\textbf{G}_{N_{\rm rx}}[m]\right)$ and $\textbf{X}[m]=\left(\textbf{I}_{N_{\rm rx}}\otimes\left(\textbf{x}^{T}[m]\otimes\textbf{I}_{M}\right)\right)$.
We employ two maximum a posteriori ratio test (MAPRT) detectors with two levels of complexity: the clutter-unaware detector and the clutter-aware detector, proposed in [12] and [40, Lem. 2], respectively. The clutter-unaware detector ignores the presence of clutter for the sake of complexity reduction, whereas the clutter-aware detector accounts for unknown clutter. The corresponding test statistics are given by
$T_{\text{c-unaware}}=\textbf{a}^{H}\textbf{C}^{-1}\textbf{a},$ (11)
$T_{\text{c-aware}}=\begin{bmatrix}\textbf{a}\\ \textbf{b}\end{bmatrix}^{H}\left(\begin{bmatrix}\textbf{C}&\textbf{E}\\ \textbf{E}^{H}&\textbf{D}\end{bmatrix}^{-1}-\begin{bmatrix}\textbf{0}&\textbf{0}\\ \textbf{0}&\textbf{D}^{-1}\end{bmatrix}\right)\begin{bmatrix}\textbf{a}\\ \textbf{b}\end{bmatrix},$ (12)
respectively, where
$\textbf{a}=\sum_{m=1}^{L_{d}}\textbf{G}^{H}[m]\textbf{y}[m],\quad\textbf{b}=\sum_{m=1}^{L_{d}}\textbf{X}^{H}[m]\textbf{y}[m],$ (13)
$\textbf{C}=\sum_{m=1}^{L_{d}}\textbf{G}^{H}[m]\textbf{G}[m]+\sigma_{n}^{2}\textbf{I}_{N_{\rm rx}N_{\rm tx}},$ (14)
$\textbf{D}=\sum_{m=1}^{L_{d}}\textbf{X}^{H}[m]\textbf{X}[m]+\sigma_{n}^{2}\textbf{R}^{-1},$ (15)
$\textbf{E}=\sum_{m=1}^{L_{d}}\textbf{G}^{H}[m]\textbf{X}[m].$ (16)
III Reliability and Delay Analysis for URLLC
For most URLLC applications, short codewords are usually needed to satisfy latency constraints where codes with short blocklengths, e.g., 50-400 symbols are employed. Short blocklengths result in decoding errors; therefore, to ensure reliability, it is essential to consider the finite blocklength regime for modeling the decoding error probability (DEP) [41, 42, 8].
We consider a maximum DEP threshold, representing the reliability requirement, together with a maximum transmission delay threshold as the URLLC requirements. In this section, we derive an upper bound on the DEP and the transmission delay, both crucial aspects considered as URLLC requirements.
In the finite blocklength regime, the communication data cannot be transmitted without error. From [9], ergodic data rate of UE $i$ can be approximated as
$R_{i}\!\approx\!\mathbb{E}\Bigg{\{}\!\!\left(1\!-\!\beta\right)\!\log_{2}{\left(1+\mathsf{SINR}^{\rm(dl)}_{i}\right)}\!-\!\frac{Q^{-1}(\epsilon_{i})}{\ln(2)}\!\sqrt{\!\frac{(1\!-\!\beta)V_{i}}{L}}\Bigg{\}},\vskip 0.0pt$ (17)
where $\beta=\frac{L_{p}}{L}$, $\epsilon_{i}$ denotes the DEP when transmitting $b_{i}$ bits to UE $i$, $\mathsf{SINR}^{\rm(dl)}_{i}$ is the instantaneous downlink communication SINR for UE $i$, $V_{i}=1-\left(1+\mathsf{SINR}^{\rm(dl)}_{i}\right)^{-2}$ is the channel dispersion, and $Q(\cdot)$ refers to the Gaussian Q-function. Due to the fact that $V_{i}<1$, the ergodic data rate can be lower bounded by
$R_{i}$ $\!\geq\!\left(1\!-\!\beta\right)\!\mathbb{E}\Bigg{\{}\!\!\log_{2}{\left(1+\mathsf{SINR}^{(\rm dl)}_{i}\right)}\!\!\Bigg{\}}\!-\!\frac{Q^{-1}(\epsilon_{i})}{\ln(2)}\sqrt{\frac{(1-\beta)}{L}}.$ (18)
Moreover, given that only $\mathbb{E}\left\{\textbf{h}_{i}^{H}\textbf{w}_{i}\right\}$ is known at UE $i$, and according to [34, Thm. 6.1] and [40, Lem. 1],
$\mathbb{E}\left\{\log_{2}{\left(1+\mathsf{SINR}^{\rm(dl)}_{i}\right)}\right\}\geq\log_{2}{\left(1+\overline{\mathsf{SINR}}^{(\rm dl)}_{i}\right)}$ (19)
where
$\overline{\mathsf{SINR}}^{(\rm dl)}_{i}=\frac{\rho_{i}\mathsf{b}_{i}^{2}}{\sum_{j=0}^{{\rm N_{ue}}}\rho_{j}\mathsf{a}_{i,j}^{2}+\sigma_{n}^{2}},\quad i=1,\ldots,N_{\rm ue}$ (20)
with
$\mathsf{b}_{i}=\left|\mathbb{E}\left\{\textbf{h}_{i}^{H}\textbf{w}_{i}\right\}\right|,\quad\mathsf{a}_{i,i}=\sqrt{\mathbb{E}\left\{\left|\textbf{h}_{i}^{H}\textbf{w}_{i}\right|^{2}\right\}-\mathsf{b}_{i}^{2}}$ (21)
$\mathsf{a}_{i,j}=\sqrt{\mathbb{E}\left\{\left|\textbf{h}_{i}^{H}\textbf{w}_{j}\right|^{2}\right\}},\quad j=0,1,\ldots,N_{\rm ue},\quad j\neq i.$ (22)
The expectations are taken with respect to the random channel realizations. Now, using (19) and substituting $R_{i}=\frac{b_{i}}{L}$ into (18), we obtain an upper bound for the DEP as
$\epsilon_{i}\leq\epsilon_{i}^{\rm(ub)}\!\triangleq\!Q\!\left(\sqrt{L-L_{p}}\left[\ln\left(1+\overline{\mathsf{SINR}}^{(\rm dl)}_{i}\right)-\frac{b_{i}\ln{2}}{L-L_{p}}\right]\right).$ (23)
In this paper, we focus on the transmission delay and leave the analysis of E2E delay as future work. Let $D_{i}^{t}$ denote the transmission delay of UE $i$, expressed as
$D_{i}^{\rm t}=\frac{T}{1-\epsilon_{i}}=\frac{L}{B(1-\epsilon_{i})},\quad i=1,\ldots,N_{\rm ue}$ (24)
where $T=\frac{L}{B}$ is the time duration of one URLLC transmission with blocklength $L$. To satisfy the reliability requirement, $\epsilon_{i}$ should be less than the maximum tolerable DEP denoted by $\epsilon_{i}^{(\rm th)}$. Then, since $\epsilon_{i}\leq\epsilon_{i}^{\rm(ub)}\leq\epsilon_{i}^{\rm(th)}$, the transmission delay is upper-bounded as
$D_{i}^{\rm t}\leq\frac{L}{B\left(1-\epsilon_{i}^{\rm(th)}\right)}\triangleq D_{i}^{\rm(ub)}\leq D_{i}^{\rm(th)},$ (25)
where $D_{i}^{\rm(th)}$ is the maximum tolerable delay by UE $i$ and $D_{i}^{\rm(ub)}\leq D_{i}^{\rm(th)}$ should be satisfied to guarantee the delay requirement. This implies that the blocklength cannot exceed $D^{\rm(th)}_{i}B(1-\epsilon_{i}^{\rm(th)})$. Thus, we can define the maximum tolerable blocklength by $L_{\rm max,c}$, where
$L_{\rm max,c}=\min\left\{D^{\rm(th)}_{i}B\left(1-\epsilon_{i}^{\rm(th)}\right)|i=1,\ldots,N_{\rm ue}\right\}.$ (26)
IV Sensing Analysis
We assess the sensing performance in terms of detection probability under a certain false alarm probability and refreshing rate in terms of the number of sensing measurement updates per time unit. Detection probability refers to the likelihood of correctly identifying the presence of a target, whereas false alarm probability denotes the likelihood of incorrectly detecting a target when none is present.
Maximizing the sensing SINR enhances the probability of target detection under a fixed false alarm probability [7, Chapters 3 and 15]. For other sensing tasks—such as estimating range, velocity, and similar parameters— it is also desired to maintain the sensing SINR above a required threshold denoted by $\gamma_{\rm s}$, i.e., $\mathsf{SINR}_{\rm s}\geq\gamma_{\rm s}$.This requirement serves as a key motivation for the optimization problem formulated in Section  VI. In this work, we consider a minimum sensing SINR and refreshing rate as the sensing requirements.
The sensing SINR is a function of symbols, which vary randomly for different blocklength values. This implies that although the symbols are known when processing the received sensing signals, they cannot be known during the resource allocation phase. To this end, we take the expectation with respect to the random symbols. Thus, the average sensing SINR would be
$\overline{\mathsf{SINR}}_{\rm s}$ $=\frac{M\bm{\rho}^{T}\textbf{A}_{D}\bm{\rho}}{MN_{\rm rx}\sigma_{n}^{2}+\bm{\rho}^{T}\textbf{B}_{D}\bm{\rho}},$ (27)
where $\textbf{A}_{D}$ and $\textbf{B}_{D}$ are diagonal matrices with
$\left[\textbf{A}_{D}\right]_{ii}$ $=\sum_{r=1}^{N_{\rm rx}}\sum_{k=1}^{N_{\rm tx}}\beta_{r,k}\left(\textbf{W}_{k}^{H}\textbf{a}^{*}(\varphi_{k},\vartheta_{k})\textbf{a}^{T}(\varphi_{k},\vartheta_{k})\textbf{W}_{k}\right)_{ii},$ (28)
$\left[\textbf{B}_{D}\right]_{ii}$ $=\sum_{r=1}^{N_{\rm rx}}\sum_{k=1}^{N_{\rm tx}}\mathrm{tr}\left(\textbf{R}_{\mathrm{rx},(r,k)}\right)\left(\textbf{W}_{k}^{H}\textbf{R}^{T}_{{\rm tx},(r,k)}\textbf{W}_{k}\right)_{ii}.$ (29)
According to the 3GPP specification [6], refreshing rate is defined as the rate at which the sensing result is generated by the sensing system. It is the inverse of the time elapsed between two successive sensing results. In our system model, the sensing information is updated once throughout each transmission block. Thus, the sensing service latency is $\frac{L}{B}$ and the refreshing rate, denoted as $R_{\rm s}$, is
$R_{\rm s}=\frac{B}{L}\quad(\rm updates/second).$ (30)
V E2E Power Consumption Modeling
Compared to communication networks, ISAC networks are expected to consume more power due to sensing tasks. In general, E2E power consumption in a system with virtualized C-RAN architecture is consisting of two main components: i) the radio site power consumption, including the AP power consumption and ii) the power consumption at the edge cloud, denoted by $P_{\rm cloud}$ [17].
In the considered virtualized C-RAN architecture, all processing is done in the cloud. Let $P_{\textrm{AP},0}^{\rm tx}$ and $P_{\textrm{AP},0}^{\rm rx}$ be the static power consumption of the transmit ISAC AP and the receive sensing AP, respectively. The total power consumption, taking into account both communication and sensing, can be expressed as
$P_{\rm total}$ $=\underbrace{\Delta^{\rm tr}P_{\rm tr}+\sum_{k=1}^{N_{\rm tx}}P_{\textrm{AP},0}^{\rm tx}+\sum_{r=1}^{N_{\rm rx}}P_{\textrm{AP},0}^{\rm rx}}_{\textrm{radio site power consumption}}+P_{\rm cloud}$ (31)
where $\Delta^{\rm tr}$ is the slope of load-dependent transmission power consumption of each AP. The average power consumption for transmission is given as $P_{\rm tr}=\sum_{j=0}^{N_{\rm ue}}\rho_{j}=\|\bm{\rho}\|^{2}$, where $\bm{\rho}=[\sqrt{\rho_{0}}\ \ldots\ \sqrt{\rho_{N_{\rm ue}}}]^{T}$ due to the unit-power centralized precoding vectors. The power consumption at the cloud is modeled as
$P_{\rm cloud}$ $=P_{\rm fixed}+\frac{1}{\sigma_{\rm cool}}\Bigg{(}N_{\rm GPP}P_{\rm cloud,0}^{\rm proc}+\Delta_{\rm cloud}^{\rm proc}\frac{C_{\rm cloud}}{C_{\rm max}}\Bigg{)},$ (32)
where $P_{\rm fixed}$ is the fixed power consumption at the cloud, which is independent of the load. $N_{\rm GPP}=\lceil\frac{C_{\rm cloud}}{C_{\rm max}}\rceil$ is the number of required general-purpose processors (GPPs), and $P_{\rm cloud,0}^{\rm proc}$ is the processing power consumption in the idle mode. $\sigma_{\rm cool}\in(0,1]$ and $\Delta_{\rm cloud}^{\rm proc}$ denote the cooling efficiency of the cloud and the slope of the load-dependent power consumption for processing at the digital unit (DU) in the cloud, respectively. Moreover, $C_{\rm max}$ and $C_{\rm cloud}$ are the maximum processing capacity of the processing resources in each GPP and the total processing resource utilization in GOPS, respectively [18, 43]. The processing resource utilization can be expressed as
$C_{\rm cloud}=C_{\rm proc}^{\rm c}+C_{\rm proc}^{\rm s}$ (33)
where $C_{\rm proc}^{\rm c}$ and $C_{\rm proc}^{\rm s}$ are the processing resource utilization due to communication and sensing tasks, respectively [^2] In this paper, we focus on the GOPS analysis by taking into account only physical-layer communication and sensing processing and neglect high-layer operations..
In the following parts, the required GOPS for communication and sensing in our system model is computed, respectively. We assess the computational complexity where only the numbers of real multiplications and divisions are counted. Each complex multiplication is equal to four real multiplications. We also consider memory overhead in arithmetic operation calculations by multiplying each operation by two as done in [44, 18]. Hence, each complex multiplication is counted as 42=84$\cdot$ 2=84 ⋅ 2 = 8 operations in computing the total GOPS.
V-A GOPS Analysis of Digital Operations for Communication at the Cloud
In this section, we analyze the GOPS for digital signal processing corresponding to the communication tasks including the uplink channel estimation and downlink transmission. To compute the number of real multiplications, we mainly follow the GOPS analysis in [37, App. B], [17].
Let $C_{\textrm{ch-est}}$ denote the computational complexity of the LMMSE channel estimation approach for all the APs. To compute the channel estimates, we first obtain the vectors $\textbf{y}_{t_{i},k}^{p}$. From [37, App. B], the multiplication of one matrix of size $M\times L_{p}$ with a vector of size $L_{p}\times 1$ results in $ML_{p}$ complex multiplications. Hence, obtaining $\textbf{y}_{t_{i},k}^{p}$ for $N_{\rm ue}$ UEs at all $N_{\rm tx}$ APs costs $8ML_{p}N_{\rm ue}N_{\rm tx}$ real multiplications/division in total, if $L_{p}\geq N_{\rm ue}$. Otherwise, the number of real multiplications/divisions would be $8ML_{p}^{2}N_{\rm tx}$. Moreover, we need to compute the matrices $\textbf{R}_{i,k}^{\prime}$ and $\bm{\Psi}_{t_{i},k}^{\prime}$, given that $\bar{\textbf{h}}_{i,k}$ and $\textbf{R}_{i,k}$ are known. However, this pre-computation can be neglected since the channel statistics are usually constant for a while and there is no need to compute them every coherence block. We use the LMMSE channel estimation approach in [38] which costs $8M^{2}N_{\rm ue}N_{\rm tx}$ real operations. Then, $C_{\textrm{ch-est}}$ is equal to
$C_{\textrm{ch-est}}=\left\{\begin{matrix}(8ML_{p}+8M^{2})N_{\rm ue}N_{\rm tx},&L_{p}\geq N_{\rm ue}\\ 8ML_{p}^{2}N_{\rm tx}+8M^{2}N_{\rm ue}N_{\rm tx},&L_{p}<N_{\rm ue}.&\end{matrix}\right.$ (34)
The number of real multiplications/divisions to compute centralized RZF precoding vector for all the UEs from [37, App. B] is
$C_{\textrm{prec-comp}}$ $=\underbrace{\frac{8\left((MN_{\rm tx})^{2}+MN_{\rm tx}\right)}{2}N_{\rm ue}}_{\textrm{inside the parentheses}}$
$\quad+\underbrace{\frac{8\left((MN_{\rm tx})^{3}-MN_{\rm tx}\right)}{3}}_{\rm inversion}+N_{\rm ue}\cdot\underbrace{8(MN_{\rm tx})^{2}}_{\textrm{for}(.)^{-1}\hat{\textbf{h}}_{i}}$
$\quad+N_{\rm ue}\cdot\left(\underbrace{8MN_{\rm tx}}_{\textrm{computing }||\textbf{w}_{i}||}+\underbrace{4MN_{\rm tx}}_{\textrm{normalization}}\right)$
$=\left(12(MN_{\rm tx})^{2}+16MN_{\rm tx}\right)N_{\rm ue}+\frac{8\left((MN_{\rm tx})^{3}-MN_{\rm tx}\right)}{3}.$ (35)
Reciprocity calibration and multiplication of the symbols by the precoding vectors, each costs $8L_{d}MN_{\rm ue}N_{\rm tx}$ real operations [45, 18]. Multiplying by the power coefficients also costs $4L_{d}MN_{\rm ue}N_{\rm tx}$. Finally, the GOPS corresponding to communication processing (i.e., channel estimation, precoding and reciprocity calibration) is computed as in (36) where we divided the total giga operations by the coherence time $L/B$.
$C_{\textrm{proc}}^{\rm c}=\frac{B}{L10^{9}}\Bigg{(}C_{\textrm{ch-est}}+C_{\textrm{prec-comp}}+20L_{d}MN_{\rm ue}N_{\rm tx}\Bigg{)}.$ (36)
TABLE II: COMPUTATIONAL COMPLEXITY FOR CLUTTER-AWARE DETECTOR
Operation Computational complexity
$\textbf{G}[m],\quad\forall m$ $20\,L_{d}\,MN_{\rm tx}N_{\rm rx}$
a $8\,L_{d}\,M^{2}\,N_{\rm rx}\,N_{\rm tx}\,$
b $8\,L_{d}\,M^{2}\,N_{\rm rx}\,N_{\rm tx}$
C $4\,L_{d}MN_{\rm rx}\left(N_{\rm tx}^{2}+N_{\rm tx}\right)$
D $4L_{d}\,\left((MN_{\rm tx})^{2}+MN_{\rm tx}\right)$
E $8\,L_{d}\,M^{2}\,N_{\rm rx}\,N_{\rm tx}^{2}$
$\begin{bmatrix}\textbf{C}&\textbf{E}\\ \textbf{E}^{H}&\textbf{D}\end{bmatrix}^{-1}$
$\frac{8}{3}\left(\left((1+M^{2})N_{\rm tx}N_{\rm rx}\right)^{3}-(1+M^{2})N_{\rm tx}N_{\rm rx}\right)$
$\textbf{D}^{-1}$ $\frac{8}{3}\left((M^{2}N_{\rm tx}N_{\rm rx})^{3}-M^{2}N_{\rm tx}N_{\rm rx}\right)$
$\begin{bmatrix}\textbf{a}^{H}\!&\textbf{b}^{H}\end{bmatrix}\!\begin{bmatrix}(.)&(.)\\ (.)&(.)\end{bmatrix}\begin{bmatrix}\textbf{a}\\ \textbf{b}\end{bmatrix}$
$8\left((1+M^{2})N_{\rm tx}N_{\rm rx}\right)^{2}+8(1+M^{2})N_{\rm tx}N_{\rm rx}$
V-B GOPS Analysis of Digital Operations for Sensing at the Cloud
In this subsection, we present the GOPS analysis for sensing operations, including signal transmission and processing the received signal. The GOPS corresponding to the sensing transmissions consists of computing the sensing precoding vector $\textbf{w}_{0}$, denoted by $C_{\rm prec-comp}^{\rm s}$, and obtaining the sensing signal by multiplying the sensing symbols by the sensing precoding vector and the sensing power coefficient. Multiplying the sensing symbols by the sensing precoding vector and the sensing power coefficient costs $12L_{d}MN_{\rm tx}$ real multiplications/divisions. The unitary matrix U in (6) (subspace spanned by the UE channel estimation vectors) is already obtained when computing the RZF precoding vectors by matrix inversion and the corresponding $\mathbf{L}\mathbf{D}\mathbf{L}^{H}$ decomposition [37, Lem. B.2]. Therefore, the computational complexity of the ZF precoding vector $\textbf{w}_{0}$ is given by
$C_{\rm prec-comp}^{\rm s}$ $=8(MN_{\rm tx})^{2}+12MN_{\rm tx},$ (37)
where the first term stands for matrix-vector multiplication and the second term corresponds to the cost of computing $||\textbf{w}_{0}||$ and normalization, which are counted as $8MN_{\rm tx}+4MN_{\rm tx}$.
After transmitting the signal in downlink, the reflected signals along with the interference signals are received at the receiver APs and sent to the cloud. At the cloud, these signals are processed for a specific sensing application. In this work, we consider target detection and assess the number of real multiplications/divisions required to compute the test statistics. For target detection problems, we usually compute the test statistics and compare them with a threshold. The target is declared detected if the value of the test statistics is greater than the threshold. We assume that the threshold is constant. Therefore, we can neglect the computational complexity of obtaining the threshold. However, test statistics should be obtained for each transmission. Hence, the sensing GOPS is obtained as
$C_{\rm proc}^{\rm s}$ $=\frac{B}{L10^{9}}\Bigg{(}\underbrace{12L_{d}MN_{\rm tx}}_{\textrm{precoding + power}}+C_{\textrm{prec-comp}}^{\rm s}+\underbrace{C_{\rm se-comp}}_{\textrm{target detection}}\Bigg{)},$ (38)
where $C_{\rm se-comp}$ denotes the computational complexity for computing the detector test statistics.
Given the clutter-unaware detector in (11), the computational complexity is computed as
$C_{\rm se-comp}^{\rm c-unaware}$ $=L_{d}\Bigg{(}\underbrace{20MN_{\rm tx}N_{\rm rx}}_{\rm{for}\,\textbf{ G}}+\underbrace{8\,M^{2}\,N_{\rm rx}\,N_{\rm tx}}_{\rm{for}\,\textbf{ a}}$
$\quad+\underbrace{4\,N_{\rm rx}\left(N_{\rm tx}^{2}+N_{\rm tx}\right)M}_{\rm{for}\,\textbf{ C}}\Bigg{)}+\underbrace{8\frac{(N_{\rm tx}^{3}-N_{\rm tx})N_{\rm rx}}{3}}_{\textrm{for inverse {C}}}$
$\quad+\underbrace{8\left((N_{\rm tx}N_{\rm rx})^{2}+N_{\rm tx}N_{\rm rx}\right)}_{\textrm{for final step}},$ (39)
where we used the block diagonal structure of C in getting the inverse of it. The computational complexity for each step of computing the test statistics of clutter-aware detector in (12) are listed in Table II.
VI Joint blocklength and Power Optimization
We aim to jointly optimize the blocklength and the power control coefficients to minimize the E2E energy consumption while URLLC and sensing requirements are satisfied.  [^3] In addition to considering radio and cloud processing energy consumption, it is important to take fronthaul energy consumption into account when assessing E2E energy consumption. However, in our analysis, where the primary focus is on the interplay between processing and radio resources, we treat fronthaul energy consumption as a fixed component and do not include it in our considerations. The corresponding algorithm to this optimization problem is represented by E2E JPB-SeURLLC+. The optimization problem is formulated as follows:
$\underset{\bm{\rho}\geq\textbf{0},L>L_{p}}{\textrm{minimize}}\quad$ $E_{\rm total}=\frac{L_{d}}{B}\Delta^{\rm tr}\|\bm{\rho}\|^{2}+\frac{L}{B}\left(P_{\rm FIXED}+\frac{1}{L}f_{1}+\frac{L_{d}}{L}f_{2}\right)\,$ (40a)
subject to $\epsilon^{\rm(ub)}_{i}\leq\epsilon^{\rm(th)}_{i},\quad\forall i>0$ (40b)
$L\leq L_{\rm max,c},\quad$ (40c)
$R_{\rm s}\geq R_{\rm s}^{\rm th},\quad$ (40d)
$\overline{\mathsf{SINR}}_{\rm s}\geq\gamma_{\rm s},\quad$ (40e)
$P_{k}\leq P_{\rm tx},\quad\quad k=1,\ldots,N_{\rm tx}$ (40f)
where
$P_{\rm FIXED}=\sum_{k=1}^{N_{\rm tx}}P_{\textrm{AP},0}^{\rm tx}+\sum_{r=1}^{N_{\rm rx}}P_{\textrm{AP},0}^{\rm rx}+P_{\rm fixed}+N_{\rm GPP}\frac{P_{\rm cloud,0}^{\rm proc}}{\sigma_{\rm cool}}$ (41)
$f_{1}=\frac{\Delta_{\rm cloud}^{\rm proc}}{\sigma_{\rm cool}C_{\rm max}}\frac{B}{10^{9}}\Bigg{(}C_{\textrm{ch-est}}+C_{\textrm{prec-comp}}+C_{\textrm{prec-comp}}^{\rm s}$
$\quad+C_{\rm se-detector}\Bigg{)}$ (42)
$f_{2}=\frac{\Delta_{\rm cloud}^{\rm proc}}{\sigma_{\rm cool}C_{\rm max}}\frac{B}{10^{9}}\left(20MN_{\rm ue}N_{\rm tx}+12MN_{\rm tx}+C_{{\rm se-prep}}\right),$ (43)
where $C_{\rm se-detector}$ is the summation of the terms in $C_{\rm se-comp}$ that does not have the multiplying factor of $L_{d}$ as it is included in $f_{1}$. On the other hand, $C_{\rm se-prep}$ includes the terms with $L_{d}$ appearing in $f_{2}$. Constraints (40b) and (40c) correspond to the URLLC requirements. Constraints (40d) and (40e) correspond to the sensing requirements, where $R_{s}^{\rm th}$ is the refreshing rate threshold and $\gamma_{\rm s}$ is the required sensing SINR that is selected according to the target detection performance requirement and $P_{\rm tx}$ is the maximum transmit power per AP. We define a new variable $\bar{L}$ where $L_{d}=L-L_{p}\leq\frac{1}{\bar{L}}$. Then, the objective function (40a) is equivalent to
$\mathsf{F}\triangleq\frac{1}{B}\left(LP_{\rm FIXED}+\left(L-L_{p}\right)f_{2}+\Delta^{\rm tr}\frac{\|\bm{\rho}\|^{2}}{\bar{L}}\right).$ (44)
The aforementioned problem is challenging to solve due to its non-convex nature and the high coupling of variables. In the following theorem, we present an equivalent optimization problem by introducing newly defined auxiliary variables. This allows us to obtain a more tractable optimization problem.
Theorem 1.
Consider the problem in (40) and define the collective vectors $\textbf{r}=[r_{1}\ \ldots\ r_{N_{\rm ue}}]^{T}$ and $\bm{\chi}=[\chi_{1}\ \ldots\ \chi_{N_{\rm ue}}]^{T}$. The optimal solution $\{\bm{\rho}^{\star},L^{\star}\}$ of the problem given below is also an optimal solution to (40):
$\underset{\bm{\rho},\bm{\chi},\textbf{r}\geq\textbf{0},L>L_{p},\bar{L}>0}{\textrm{minimize}}\quad\mathsf{F}$ (45a)
[…]
$\ln\left(1+\chi_{i}\right)-\chi_{i}+\frac{\rho_{i}\mathsf{b}_{i}^{2}}{r_{i}}\geq\frac{Q^{-1}\left(\epsilon_{i}^{\rm(th)}\right)}{\sqrt{L-L_{p}}}+\frac{b_{i}\ln{2}}{L-L_{p}}$ (45c)
$\left\|\begin{bmatrix}\sqrt{2}\mathsf{a}_{i,0}\sqrt{\rho_{0}}\\ \vdots\\ \sqrt{2}\mathsf{a}_{i,N_{\rm ue}}\sqrt{\rho_{N_{\rm ue}}}\\ \sqrt{2}\mathsf{b}_{i}\sqrt{\rho_{i}}\\ \sqrt{2}\sigma_{n}\\ 1+\chi_{i}\\ r_{i}\end{bmatrix}\!\right\|\leq 1+\chi_{i}+r_{i},\forall i,$ (45d)
$\bm{\rho}^{T}\left(\gamma_{\rm s}\textbf{B}_{D}-M\,\textbf{A}_{D}\,\right)\bm{\rho}\leq-\gamma_{\rm s}MN_{\rm rx}\sigma_{n}^{2}$ (45e)
$\left\|\textbf{F}_{k}\bm{\rho}\right\|\leq\sqrt{P_{\rm tx}},\quad k=1,\ldots,N_{\rm tx},$ (45f)
$L\leq\min\left(L_{\rm max,c},\frac{B}{R_{\rm s}^{\rm th}}\right),$ (45g)
where ${\bf F}_{k}=\mathrm{diag}\left(\sqrt{\mathbb{E}\left\{\|{\bf w}_{0,k}\|^{2}\right\}},\ldots,\sqrt{\mathbb{E}\left\{\|{\bf w}_{N_{\rm ue},k}\|^{2}\right\}}\right)$.
Proof.
See Appendix A. ∎
The optimization problem in (45) is still not convex due to the non-convex constraints (45b), (45c) and (45e). The terms that destroy convexity are the convex terms $1/\overline{L}$ and $\frac{\rho_{i}\mathsf{b}_{i}^{2}}{r_{i}}$ (in terms of $\bm{\rho}$ and r) on the right-hand side of (45b) and the left-hand side of (45c), respectively. To this end, we apply the concave-convex procedure (CCP) approach to (45b) and (45c), and the FPP-SCA method [46] to (45e), wherein $-\bm{\rho}^{T}M\,\textbf{A}_{D}\bm{\rho}$ is a concave function. Moreover, to avoid any potential infeasibility issue regarding (45e) during the initial iterations of the algorithm, we add slack variable $\chi_{0}\geq 0$ and a slack penalty $\lambda$, to the convexified problem at the initial iterations. In subsequent iterations, we set $\chi_{0}$ to zero if it is less than a threshold, denoted as $\chi_{0}\leq\epsilon_{\chi}$. Finally, the convex problem that is solved at the $c^{\rm th}$ iteration becomes
$\underset{\bm{\rho},\bm{\chi},\textbf{r}\geq\textbf{0},L>L_{p},\bar{L}>0,\chi_{0}\geq 0}{\textrm{minimize}}\quad\mathsf{F}+\lambda\chi_{0}$ (46a)
$\textrm{subject to}\quad{L-L_{p}}\leq\frac{2}{\overline{L}^{(c-1)}}-\frac{\overline{L}}{\left(\overline{L}^{(c-1)}\right)^{2}},$ (46b)
$\ln\left(1+\chi_{i}\right)-\chi_{i}+2\frac{\sqrt{\rho_{i}}^{(c-1)}\mathsf{b}_{i}^{2}\,\sqrt{\rho_{i}}}{r_{i}^{(c-1)}}-r_{i}\left(\frac{\sqrt{\rho_{i}}^{(c-1)}\mathsf{b}_{i}}{r_{i}^{(c-1)}}\right)^{2}$
$\hskip 14.22636pt\geq\frac{Q^{-1}\left(\epsilon_{i}^{\rm(th)}\right)}{\sqrt{L-L_{p}}}+\frac{b_{i}\ln{2}}{L-L_{p}}$ (46c)
$\gamma_{\rm s}\bm{\rho}^{T}\textbf{B}_{D}\bm{\rho}-2M\,\Re\left(\left(\bm{\rho}^{(c-1)}\right)^{T}\textbf{A}_{D}\bm{\rho}\right)$
$\hskip 14.22636pt\leq-\gamma_{\rm s}MN_{\rm rx}\sigma_{n}^{2}-M\left(\bm{\rho}^{(c-1)}\right)^{T}\textbf{A}_{D}\bm{\rho}^{(c-1)}+\chi_{0},$ (46d)
[…]
The steps of the FPP-SCA and CCP procedure, E2E JPB-SeURLLC+ algorithm, are outlined in Algorithm 1. We empirically observed that setting $\overline{L}^{(0)}=\frac{1}{L_{\rm max}-L_{p}}$, $\rho_{0}^{(0)}=0$ and $\sqrt{\rho_{i}}^{(0)}=10^{-3}\sqrt{P_{\rm tx}/N_{\rm ue}}$ for $i>0$ yields satisfactory results.
Algorithm 1 FPP-SCA and CCP Procedure for Solving Problem (45), E2E JPB-SeURLLC+ algorithm
1:Initialization: Initialize $\rho_{0}^{(0)}=0$, $\sqrt{\rho_{i}}^{(0)}=10^{-3}\sqrt{P_{\rm tx}/N_{\rm ue}}$ for $i>0$, $\textbf{r}^{(0)}\geq\textbf{0}$, and blocklength $\overline{L}^{(0)}=\frac{1}{L_{\rm max}-L_{p}}$. Set solution accuracy parameters $\epsilon,\epsilon_{\chi}>0$, and $\lambda>0$. Set iteration counter $c=0$, maximum number of iterations $c_{\rm max}$, the initial objective value $\mathsf{F}^{(0)}=\infty$ and define the improvement metric as $\Delta\mathsf{F}^{(c)}=\mathsf{F}^{(c-1)}-\mathsf{F}^{(c)}$, and $\Delta\mathsf{F}^{(0)}=\infty$.
[…]
3:while $\Delta\mathsf{F}^{(c-1)}\geq\epsilon$ and $c\leq c_{\rm max}$ do
4:  Solve the convexified problem in (46) using the previous iterates $\bm{\rho}^{(c-1)}$, $\overline{L}^{(c-1)}$, and $\textbf{r}^{(c-1)}$ as constants, and update $\bm{\rho}^{(c)}$, $\overline{L}^{(c)}$, and $\textbf{r}^{(c)}$ accordingly.
5:  if $\chi_{0}<\epsilon_{\chi}$ then
6:   Set $\chi_{0}=0$ for the next iteration.
7:  end if
8:  $c\leftarrow c+1$
9:end while
10:Output: Transmit power coefficients $\bm{\rho}^{(c)}$ and the optimized blocklength $L^{(c)}$.
VII Numerical Results
[…]
The large-scale fading coefficients, shadowing parameters, probability of LOS, and the Rician factors are simulated based on the 3GPP Urban Microcell model, defined in [47, Table B.1.2.1-1, Table B.1.2.1-2, Table B.1.2.2.1-4]. The path losses for the Rayleigh fading target-free channels are also modeled by the 3GPP Urban Microcell model with the difference that the channel gains are multiplied by an additional scaling parameter equal to 0.30.30.3 to suppress the known parts of the target-free channels due to LOS and permanent obstacles [40]. The sensing channel gains are computed by the bi-static radar range equation [7]. The carrier frequency, the bandwidth, and the noise variance are set to 1.91.91.9 GHz, 200200200 KHz, and 114-114- 114 dBm, respectively. The number of pilot symbols is $L_{p}=10$. The regularization parameter $\delta$ in (5) is set to the noise variance.
The spatial correlation matrices for the communication channels are generated by using the local scattering model in [34, Sec. 2.5.3]. The RCS of the target is modeled by the Swerling-I model with $\sigma_{\rm rcs}=0$ dBsm. For all the UEs, the packet size, maximum transmission delay, and DEP threshold are $b_{i}=256$ bits, $D_{i}^{\rm(th)}=1$ ms, and $\epsilon_{i}^{\rm(th)}=10^{-5}$, respectively. The sensing SINR threshold is $\gamma_{s}=0$ dB, the false alarm probability threshold is $P_{\rm fa}=0.03$, and the refreshing rate threshold is $R_{\rm s}^{\rm th}=10$ updates per second [6, Table 6.2-1], unless otherwise stated. The remaining parameters are detailed in Table III, where the values are consistent with those in [18].
TABLE III: SIMULATION PARAMETERS
$\Delta^{\rm tr},\sigma_{\rm cool}$ 4, 0.9 $P_{\rm fixed}$ 120 W
$P_{\textrm{AP},0}^{\rm tx},P_{\textrm{AP},0}^{\rm rx}$ $6.8\cdot M$ W $P_{\rm cloud,0}^{\rm proc}$ 81 W
$\Delta_{\rm cloud}^{\rm proc}$ 288 W $C_{\rm max}$ 700.94 GOPS
We compare the performance of the proposed algorithm, E2E JPB-SeURLLC+, which aims to jointly optimize power and blocklength to minimize E2E energy consumption, against two benchmark schemes: (i) JPB-SeURLLC+, which minimizes only the transmission energy consumption, and (ii) E2E JPB-URLLC, which targets E2E energy minimization in URLLC systems without sensing integration. We also evaluate performance under both clutter-aware and clutter-unaware MAPRT detectors, as defined in (11) and (12), respectively. Our analysis focuses on the total energy consumed to complete one cycle of communication and sensing tasks. The communication task requires satisfying a maximum DEP threshold for each UE, while the sensing task involves detecting a target at a candidate location with a specified false alarm probability threshold. The total E2E energy includes ISAC transmission, sensing reception, and processing components for both tasks.
Figure 3: Convergence condition for E2E JPB-SeURLLC+, JPB-SeURLLC+, and E2E JPB-URLLC algorithms with clutter-aware detector and $N_{\rm rx}=2$.
Figure 4: (a) Detection probability, (b) transmission energy, and (c) E2E energy vs. sensing SINR threshold with $N_{\rm rx}=2$.
Figure 5: Maximum blocklength threshold and network availability vs. refreshing rate threshold for $N_{\rm rx}=2$ and the E2E JPB-SeURLLC+ algorithm.
Figure 6: (a) Detection probability, (b) transmission, and (c) total energy consumption vs. DEP threshold for $N_{\rm rx}=2$, $\gamma_{\rm s}=0$ dB, and $R_{\rm s}^{\rm th}=1000$ updates/s.
Figure 7: Breakdown of energy consumption for completing one sensing and communication task with clutter-aware detector, $N_{\rm rx}=2$, and $\gamma_{\rm s}=0$ dB.
Figure 8: (a) Computational complexity for one sensing task, (b) sensing GOPS, (c) number of required GPPs vs. blocklength.
(a)(b)
Figure 9: (a) Number of antenna elements per AP, and (b) total energy consumption for a minimum detection probability of 0.90.90.9 with E2E JPB-SeURLLC+ algorithm and $N_{\rm rx}=1,2$.
Fig. 3 illustrates the convergence behavior of the proposed and benchmark algorithms. The objective of the E2E algorithms is to minimize the total energy consumption, whereas the JPB-SeURLLC+ algorithm focuses solely on minimizing transmission energy consumption. In the figure, the values for JPB-SeURLLC+ are scaled by $10^{5}$ since they are significantly lower compared to the total energy values. In the algorithms, the solution accuracy parameters are set as $\epsilon=10^{-3}$, $\epsilon_{\chi}=10^{-6}$, $\lambda=10$, and the maximum iteration is set to $c_{\rm max}=30$.
In the following, we first compare the performance of the proposed E2E JPB-SeURLLC+ algorithm with relevant benchmarks, followed by a comprehensive analysis of the proposed algorithm under various communication and sensing requirements. Then, we investigate the interplay between the number of receive APs, the number of antenna elements per AP, and the complexity level of the detector to balance the sensing performance and energy consumption in the system. Detection probability results are obtained empirically through Monte Carlo simulation.
VII-A Impact of Sensing and Communication Requirements
The effect of the sensing SINR threshold $\gamma_{\rm s}$ is evaluated in Fig. 4 for $N_{\rm rx}=2$. As shown in Fig. 4(a), detection probability generally increases with $\gamma_{s}$, but performance varies based on the detector and resource allocation strategy. Clutter-aware detectors consistently outperform clutter-unaware ones due to their advanced signal processing. As this detector estimates the target-free channels, its performance does not change significantly with increasing $\gamma_{\rm s}$ while the performance of clutter-unaware detector accordingly. The proposed E2E JPB-SeURLLC+ with a clutter-aware detector achieves the highest detection probability up to 0.980.980.98, while the same algorithm with a clutter-unaware detector reaches only detection probability of 0.630.630.63. The JPB-SeURLLC+ algorithm that prioritizes minimizing transmission energy consumption, as opposed to E2E energy minimization, tend to result in lower detection performance. This is because the E2E algorithms try to minimize the total energy consumption by reducing the blocklength. This approach allows higher transmission power levels–as shown in Fig. 4(b)– to meet the requirements, which in turn enhance detection capability. In contrast, E2E JPB-URLLC fails to exceed 0.5 detection probability, even with a clutter-aware detector—highlighting the need to jointly optimize sensing and communication in URLLC systems.
[…]
Fig. 5 illustrates the maximum blocklength threshold and network availability as a function of the refreshing rate threshold. The network availability represents the percentage of cases where the optimization problem is feasible, meaning that all the requirements can be satisfied [16]. As the required refreshing rate exceeds 1000 updates/s, the system becomes increasingly constrained, resulting in a notable reduction in allowable blocklength. Specifically, the maximum blocklength is nearly halved when $R_{\rm s}^{\rm th}=2000$. However, such reduced blocklengths may not satisfy reliability requirements, leading to infeasibility. Accordingly, the network availability drops to 80% at 1800 updates/s and further declines to 60% at 2000 updates/s.
Figs. 6a–c illustrate the impact of the DEP threshold on sensing performance, transmission energy, and total energy consumption, respectively. As shown in Fig. 6(a), stricter reliability requirements slightly enhance detection probability, since the communication task requires higher power and/or longer blocklengths to meet lower DEP thresholds, which in turn enhances the sensing performance. However, this improvement comes at the cost of increased energy consumption, as evident in Fig. 6(b) and Fig. 6(c). Notably, the E2E JPB-SeURLLC+ algorithm consistently outperforms the JPB-SeURLLC+ algorithm in terms of both sensing performance and energy consumption.
It is worth mentioning that higher delay threshold allows the system to operate with higher blocklength. However, it does not effect the results since the algorithm still chooses a smaller blocklength to minimize the energy consumption.
VII-B Impact of Detector and Number of APs
Fig. 7 presents a detailed breakdown of task-oriented energy consumption across system components and operations, including ISAC transmit APs, sensing receive APs, communication and sensing processing, and an “Others” category representing load-independent and idle-mode power consumption in the cloud. As shown, the majority of energy consumption is attributed to sensing processing tasks, ISAC transmit APs, and the cloud’s load-independent and idle-mode power usage.
To further explore the system behavior, Figs. 8a-c analyze the impact of blocklength, detector type, and the number of sensing receive APs on computational complexity, GOPS, and the number of required GPPs, respectively. This multi-dimensional analysis highlights how detector complexity and network configuration influence overall system efficiency. As shown in Fig. 8(a) and Fig. 8(b), clutter-aware detector with $N_{\rm rx}=2$ has the highest computational complexity and GOPS requirements among other configurations. Although longer blocklengths generally increase the number of operations for one task, they reduce the number of operations per unit time, thereby lowering GOPS requirements. This, in turn, reduces the number of active GPPs as shown in Fig. 8(c) and consequently decreases energy consumption through the $N_{\rm GPP}P_{\rm cloud,0}^{\rm proc}$ term in equation (32). Clutter-unaware detector maintains much lower complexity compared to the clutter-aware detector. However, a comparison between the clutter-aware and clutter-unaware curves reveals distinct trends. For the clutter-aware detector, the computational complexity remains approximately constant as blocklength increases, whereas for the clutter-unaware detector, complexity grows significantly with larger blocklengths. Interestingly, the opposite trend is observed in the GOPS curves. This behavior stems from the interplay between blocklength-dependent and blocklength-independent terms in the computational complexity expressions—specifically, equation (V-B) for the clutter-unaware detector and Table II for the clutter-aware detector. Based on these observations, we conclude that for the clutter-aware detector, blocklength-independent terms dominate the computational complexity, whereas for the clutter-unaware detector, blocklength-dependent terms are the primary contributors.
[…]
VIII Conclusion and Future Directions
[…]
Appendix A Proof of Theorem 1
Let us define a new optimization variable, denoted by $\overline{L}$, where $L-L_{p}\leq 1/\overline{L}$. Then, minimizing the objective function is equivalent to minimizing the convex function (quadratic-over-linear plus affine function) $\mathsf{F}$ in (44). Minimizing this function, at the optimal solution, leads to $\overline{L}=1/L_{d}=1/(L-L_{p})$. The reliability constraints in (40b) can be written in the form of
$\ln\left(1+\frac{\rho_{i}\mathsf{b}_{i}^{2}}{\sum_{j=0}^{{\rm N_{ue}}}\rho_{j}\mathsf{a}_{i,j}^{2}+\sigma_{n}^{2}}\right)\!\geq\frac{Q^{-1}\left(\epsilon_{i}^{(\rm th)}\right)}{\sqrt{L-L_{p}}}+\frac{b_{i}\ln{2}}{L-L_{p}}$ (47)
according to (23), where $\overline{\mathsf{SINR}}^{\rm(dl)}_{i}$ is substituted by (20). To handle the non-convexity of the left-hand side in (47), we define a new variable $\chi_{i}$ and use fractional programming [48] to write the left-hand side as
$\ln(1+\chi_{i})-\chi_{i}+(1+\chi_{i})\frac{\rho_{i}\mathsf{b}_{i}^{2}}{\sum_{j=0}^{{\rm N_{ue}}}\rho_{j}\mathsf{a}_{i,j}^{2}+\rho_{i}\mathsf{b}_{i}^{2}+\sigma_{n}^{2}}.$ (48)
Moreover, to represent the upper bound to $\left(\sum_{j=0}^{{\rm N_{ue}}}\rho_{j}\mathsf{a}_{i,j}^{2}+\rho_{i}\mathsf{b}_{i}^{2}+\sigma_{n}^{2}\right)/(1+\chi_{i})$, we introduce the optimization variable $r_{i}$, similarly as in [18], for $i=1,\ldots,N_{\rm ue}$, where
$\frac{\sum_{j=0}^{{\rm N_{ue}}}\rho_{j}\mathsf{a}_{i,j}^{2}+\rho_{i}\mathsf{b}_{i}^{2}+\sigma_{n}^{2}}{1+\chi_{i}}\leq r_{i},$ (49)
which can be written as a second-order cone (SOC) constraint in (45d). We then re-cast the constraint in (48) as
$\ln\left(1+\chi_{i}\right)-\chi_{i}+\frac{\rho_{i}\mathsf{b}_{i}^{2}}{r_{i}}\geq\frac{Q^{-1}\left(\epsilon_{i}^{\rm(th)}\right)}{\sqrt{L-L_{p}}}+\frac{b_{i}\ln{2}}{L-L_{p}}$ (50)
which will not destroy optimality since we want to minimize $r_{i}$ to increase the left-hand side of the SINR constraint.
From (27), the sensing constraint in (40e) are expressed as
$\bm{\rho}^{T}\left(\gamma_{\rm s}\textbf{B}_{D}-M\,\textbf{A}_{D}\,\right)\bm{\rho}\!\leq\!-\gamma_{\rm s}MN_{\rm rx}\sigma_{n}^{2}.$ (51)
Finally, we define $\textbf{F}_{k}=\textrm{diag}\left(\sqrt{\mathbb{E}\left\{\|\textbf{w}_{0,k}\|^{2}\right\}},\ldots,\sqrt{\mathbb{E}\left\{\|\textbf{w}_{N_{\rm ue},k}\|^{2}\right\}}\right)$ and rewrite the per-AP power constraints in (40f) in SOC form in terms of $\bm{\rho}$ as $,\left\|\textbf{F}_{k}\bm{\rho}\right\|\leq\sqrt{P_{\rm tx}},$ for $k=1,\ldots,N_{\rm tx}$.